A particle with mass 1.81 x 10-3 kg and a charge of 1.22 x 10-8 C has, at a given instant, a velocity 𝑣𝑣⃑ = (3.00 Γ— 104 π‘šπ‘š 𝑠𝑠 )πš₯πš₯Μ‚. What are the magnitude and direction of the particle’s acceleration produced by a uniform magnetic field 𝐡𝐡�⃑ = (1.63 𝑇𝑇)πš€πš€Μ‚+ (0.980 𝑇𝑇)πš₯πš₯Μ‚Μ‚?


1
Expert's answer
2022-03-23T13:52:52-0400

The magnetic field force

"\vec F=q\vec v\times \vec B"

"\vec F=1.22*10^{-8}*(3.00\times 10^4\hat j)\times (1.63\hat i+0.980\hat j)\\
=-5.97*10^{-4}\hat k"

The acceleration

"\vec a=\frac{\vec F}{m}=\frac{-5.97*10^{-4}\hat k}{1.81*10^{-3}}=-0.33\hat k\:\rm m/s^2"